OpenStudy (mirracle_scott):

The sum of Sally and Beth's age is 60. Six years ago, Sally was 3 times as old as Beth. What are their present ages?

OpenStudy (mariamikayla):

Let it be S-the age of Sally and B-the age of Beth S+B=60 S-6=3(B-6) S-6=3B-18 S=3B-18+6 S=3B-12 3B-12+B=60 4B=60+12 4B=72 B=72/4 B=18 B=60-18 B=42

OpenStudy (danettsep):

and then you just fill it out :)